A car travelling at a uniform speed of 120kmh passes two stations in 4mins, calculate the distance between the two stations


1
Expert's answer
2021-05-07T06:40:57-0400

We can find the distance between the two stations from the formula:


"d=vt,""d=120\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}\cdot4\ s= 133\ m."